OpenStudy (anonymous):

In June, Katherine bought 21 cans of food and 3 bags of litter for a total of $42. In July she spent $60 on 32 cans of food and 4 bags of litter. What is the cost of one bag of litter?

OpenStudy (anonymous):

In June, Katherine bought 21 cans of food and 3 bags of litter for a total of $42. In July she spent $60 on 32 cans of food and 4 bags of litter. What is the cost of one bag of litter?

OpenStudy (anonymous):

let, the cost of one bag of litter is $x and one can of food $y then, 21y + 3x = 42.........(1) and 32y + 4x = 60...........(2) solving (1) and (2) x= 7 and y = 1 so, the cost of one bag of litter is $7